# LoRA Learns Less and Forgets Less

### v1

@@ 6 Discussion
Does the difference between LoRA and full finetuning decrease with model size?
Studies in the past have hinted at a relationship between the effectiveness of finetuning and model size (Aghajanyan et al., 2020; Hu et al., 2021; Zhuo et al., 2024). While recent studies have successfully applied LoRA to 70B parameter models (Ivison et al., 2023; Yu et al., 2023), we leave a rigorous study of these intriguing scaling properties to future work.
@@
The observation that full finetuning tends to find high rank solutions does not rule out the possibility of low-rank solutions; rather, it shows that they are not typically found. An alternative interpretation is that the rank needed to reconstruct the weight matrix is higher than the rank needed for a downstream task.
Why does LoRA perform well on math and not code?
One hypothesis is that math datasets involve a smaller domain shift; they include a larger percentage of English and lead to decreased forgetting. The second hypothesis is that the GSM8K evaluation is too easy and does not capture the new college-level math learned in finetuning.
@@ 7 Conclusion
This work sheds light on the downstream performance of contemporary LLMs (with 7 and 13 billion parameters) trained with LoRA. Different from most prior work, we use domain-specific datasets in code and math, associated with sensitive evaluation metrics. We show that LoRA underperforms full finetuning in both domains. We also show that LoRA keeps the finetuned model’s behavior close to that of the base model, with diminished source-domain forgetting and more diverse generations at inference time. We investigate LoRA’s regularization properties, and show that full finetuning finds weight perturbations are far from being low-rank. We conclude by analyzing LoRA’s increased sensitivity to hyperparameters and highlighting best practices.
@@ Appendix A Experimental Setup
Code CPT.
  • Optimizer: Decoupled LionW ($\beta=0.9,0.95$, weight decay$=1e-6$)
  • Learning rate = $51e-$ 5 for full finetuning and LoRA after a learning rate sweep.
  • LR Scheduler: Inverse square root with warmup $t_{warmup}=500$ batches, $t_{scale}=500$ batches, $t_{cooldown}=5200$ batches $\alpha_{fdecay}=1.0$ $\alpha_{fcooldown}=0.0$
@@
  • We train the model once for 20B tokens. Then perform individual cooldowns as follows. We set a target max training duration, and define the last 20% of max training duration as the cooldown period. We retrain from the latest available checkpoint prior to the cooldown period.
  • LoRA configuration: $\alpha=32$, dropout=0.05.
  • maximum sequence length = 4096.
Math CPT.
  • Optimizer: Decoupled LionW ($\beta=0.9,0.95$, weight decay$=1e-6$).
  • Learning rate = $54e-$ 5 for LoRA and $65e-$ 6 for full finetuning.
  • We ran individual runs with the Cosine with Warmup scheduler ($\alpha_{f}=0,t_{warmup}=100$ batches), without learning rate cooldowns
  • LoRA configuration $\alpha=2r$, dropout=0.05.
  • maximum sequence length = 4096
Code IFT.
  • Optimizer: Decoupled LionW ($\beta=0.9,0.95$, weight decay$=0$)
  • LR scheduler: Cosine with Warmup scheduler with $\alpha_{f}=0.01$ and $t_{warmup}=0.1$ of the total duration.
  • LoRA configuration: $\alpha=32$, dropout=0.05.
  • maximum sequence length = 4096
Math IFT. Same as code IFT, except that
  • maximum sequence length = 1024
We compared the two optimizers by training for two epochs of Magicoder-Evol-Instruct-110K using different learning rates. We found that Decoupled LionW outperformed DecoupledAdamW on HumanEval for both LoRA and full finetuning, and across learning rates, as seen in Fig. S1.
@@ Appendix B Learning rate searches
For IFT we find that LoRA LRs should be an order of magnitude higher. For the longer CPT, these effects are more subtle.
@@ B.1 Learning rate sensitivity analysis across optimizers
Figure S1: Comparing LionW to AdamW across learning rates for two epochs of the Megicoder-Evol-Instruct-110K dataset. Left: HumanEval; right: Average of “Language Understanding” benchmarks. Both methods peak at the learning rate used in the original paper (Wei et al., 2023)
Figure S2: Sample-efficiency curves matching  Fig. 2, with all individual LoRA configurations.
Figure S3: Pareto curves for continued pretraining of Llama-2-13B on up to 20B tokens of the Starcoder-Python (Code CPT).
(a) Code CPT: Individual forgetting plots for Llama-2-7B on Starcoder-Python.
(b) Code IFT: Individual forgetting plots for Llama-2-7B on Magicoder-Evol-Instruct-110K.
(c) Math CPT: Individual forgetting plots for Llama-2-7B on OpenWebMath.
(d) Math IFT: Individual forgetting plots for Llama-2-7B on MetaMathQA.
Figure S4: Same data as Fig. 3
(a) Code CPT: Individual Pareto curves for Llama-2-7B on Starcoder-Python.
(b) Code IFT: Individual Pareto curves for Llama-2-7B on Magicoder-Evol-Instruct-110K.
(c) Math CPT: Individual Pareto curves for Llama-2-7B on OpenWebMath.
(d) Math IFT: Individual Pareto curves for Llama-2-7B on MetaMathQA.
Figure S5: Same data as in Fig. 4 plotted for individual tasks HellaSwag, ARC-Challenge and WinoGrande
@@
(a) Spectrum of $A$, $A+cB$ and $cB$ for $c=0.1$.
@@
Figure S7: Analyzing the sprectra of the sum of two $1000\times 1000$ Gaussian i.i.d matrices. $A$ and $B$ are $1000\times 1000$ random matrices with i.i.d. standard normal Gaussian entries.
Appendix C Training Datasets
C.1 MetaMathQA (Math IFT)

### v2

@@ 6 Discussion
Does the difference between LoRA and full finetuning change with model size?
Studies in the past have hinted at a relationship between the effectiveness of finetuning and model size (Aghajanyan et al., 2020; Hu et al., 2021; Zhuo et al., 2024). While recent studies have successfully applied LoRA to 70B parameter models (Ivison et al., 2023; Yu et al., 2023; Niederfahrenhorst et al., 2023; Turgutlu, 2024), and previous work shows that techniques like prompt tuning become more effective for larger models (Vu et al., 2022), we leave a rigorous study of these intriguing scaling properties to future work.
@@
The observation that full finetuning tends to find high rank solutions does not rule out the possibility of low-rank solutions; rather, it shows that they are not typically found. An alternative interpretation is that the rank needed to reconstruct the weight matrix is higher than the rank needed for a downstream task. We also only presented SVD analysis for the continued pretraining setting. It is possible that a similar analysis for the instruction finetuning setting would reveal that the full finetuning does not tend to be as high rank.
@@ 7 Conclusion
This work sheds light on the downstream performance of 7 billion parameter LLMs trained with LoRA and full finetuning. Unlike most prior work, we use domain-specific datasets in code and math, associated with sensitive evaluation metrics. We show that LoRA, with commonly used low-rank settings, underperforms full finetuning across domains. We also show that LoRA keeps the finetuned model’s behavior close to that of the base model, with diminished source-domain forgetting and more diverse generations at inference time. We show that LoRA mitigates forgetting more than classical regularization techniques, and also show that full finetuning finds weight perturbations that are far from being low-rank. We conclude by analyzing LoRA’s increased sensitivity to hyperparameters and highlighting best practices.
Acknowledgements
We would like to thank the editor and the three anonymous reviewers who provided high-quality feedback on this work. We are also grateful to Daniel Han and Damjan Kalajdzievski for carefully reading our work and pointing out the importance of setting $\alpha=2r$ for training with high ranks.
Author Contributions
D.B. led this project by developing code, running experiments, analyzing results, and writing the manuscript. J.P. ran experiments and assisted in the writing of the manuscript. J.G.O. wrote code and ran experiments. P.G. advised the SVD analysis, C.J. ran experiments, and D.K. wrote code. M.P., S.H., V.C., J.F., C.B., and J.P.C. advised this work.
@@ Appendix A Experimental Setup
LoRA configuration for all experiments.
All experiments were done with the Databricks MosaicML composer, streaming and llm-foundry libraries in conjunction with the HuggingFace peft library on 32×32$\times$32 ×H100-80GB GPUs. All experiments in the main text used the LionW optimizer (Chen et al., 2021) instead of the AdamW optimizer.
We targeted all trainable modules inside each of the $L$ Llama transformer blocks: $\{W_{q}^{(l)},W_{k}^{(l)},W_{v}^{(l)},W_{o}^{(l)},W_{\text{gate}}^{(l)},W_{\text{up}}^{(l)},W_{\text{down}}^{(l)}\}_{l=1}^{L}$. We used ranks of $r=16,64,256$ and set $\alpha=2r$, to achieve a constant scaling factor $\gamma_{r}=2$ across ranks. We use lora_dropout=0.05.
@@
For both the Code CPT and Math CPT settings, we train the model once for 20B tokens. We then perform individual cooldowns using intermediate checkpoints as follows: We set a target max training duration (e.g. 8 billion tokens), and define the last 20% of max training duration as the cooldown period. We then retrain from the latest available checkpoint prior to the cooldown period.
In all four scenarios below, we use the Llama-2-7B base model meta-llama/Llama-2-7b-hf [^13] https://huggingface.co/meta-llama/Llama-2-7b-hf. For the CPT runs, we use the meta-llama/Llama-2-7b-hf tokenizer, while for the IFT runs we use the meta-llama/Llama-2-7b-chat-hf tokenizer. [^14] https://huggingface.co/meta-llama/Llama-2-7b-chat-hf
Code CPT Llama-2-7B trained on the StarCoder-Python dataset.
  • seq_len: 4096
  • optimizer: decoupled_lionw (betas=[0.9, 0.95])
  • learning_rate: 1.0e-05 for LoRA and Full Finetuning
  • scheduler: inv_sqrt_with_warmup (t_scale=1000ba, t_warmup=1000ba, t_cooldown=5086ba, alpha_f_decay=1, alpha_f_cooldown=0). We note that this ends up looking very much like a trapezoidal schedule.
  • weight_decay: 1.0e-06
  • precision: amp_bf16
  • global_train_batch_size: 192
  • device_train_microbatch_size: 6
  • gradient_clipping: norm (threshold=1)
  • num_gpus: 32
Math CPT. Llama-2-7B trained on the OpenWebMath dataset.
  • max_seq_len: 4096
  • optimizer: decoupled_lionw (betas=[0.9, 0.95])
  • learning_rate: 1.0e-05 for full finetuning, 4.0e-05 for LoRA
  • scheduler: inv_sqrt_with_warmup (t_scale=1000ba, t_warmup=1000ba, t_cooldown=5086ba, alpha_f_decay=1, alpha_f_cooldown=0). We note that this ends up looking very much like a trapezoidal schedule.
  • weight_decay: 0
  • precision: amp_bf16
  • global_train_batch_size: 192
  • device_train_microbatch_size: 6
  • gradient_clipping: norm (threshold=1)
  • num_gpus: 32
Code IFT: Finetuning Llama-2-7B on the Magicoder-Evol-Instruct-110K dataset
  • max_seq_len: 4096
  • optimizer: decoupled_lionw (betas=[0.9, 0.95])
  • learning_rate: 2e-4 for rank $r=16,64$ and 1e-4 for $r=256$ $\alpha=2r=512$ (due to instabilities/loss spikes at 2e-4)
  • scheduler: cosine_with_warmup (alpha_f=0.01, t_warmup=0.1dur)
  • weight_decay: 0
  • precision: amp_bf16
  • global_train_batch_size: 192
  • device_train_microbatch_size: 6
  • gradient_clipping: norm (threshold=1)
  • num_gpus: 32
Math IFT: Finetuning Llama-2-7B on the MetaMathQA dataset
  • seq_len: 1024
  • optimizer: decoupled_lionw (betas=[0.9, 0.95])
  • learning_rate: Full finetuning: 1e-5, LoRA: 1e-4 for $r=16,64$, 5e-5 for $r=256$ due to instabilities.
  • scheduler: cosine_with_warmup (alpha_f=0.01, t_warmup=0.1dur)
  • weight_decay: 0
  • precision: amp_bf16
  • global_train_batch_size: 768
  • device_train_microbatch_size: 24
  • gradient_clipping: norm (threshold=1)
  • num_gpus: 32
A.1 Training the input and output embedding layers.
Vanilla LoRA and other popular methods such as QLoRA (Dettmers et al., 2024) often do not train the input and output embedding layers. Recent open-source work, [^15] https://unsloth.ai/blog/contpretraining see also the following blogpost https://www.anyscale.com/blog/fine-tuning-llms-lora-or-full-parameter-an-in-depth-analysis-with-llama-2 (Niederfahrenhorst et al., 2023) on the other hand, shows that it might be beneficial to supplement LoRA with full finetuning of these two modules (additional $\approx 200M$ parameters for a 7B model). We view this approach as a hybrid of LoRA and full finetuning, and therefore leave its empirical investigation for future work. Moreover, this hybrid approach involves further hyperparameter optimization: the input and output layers require tuning their own separate learning rates, which should typically be 2-10$\times$ smaller than the LoRA learning rates (training with a single learning rate results in instabilities).
@@ Appendix B Learning rate searches
We perform a learning rate sensitivity analysis for Llama-2-7B, trained for two epochs on the code and math IFT datasets, and followed by HumanEval and GSM8K evaluation, respectively. Fig. S1 shows that LoRA improves monotonically with learning rate up to a value at which training diverges, with best learning rates of $5e^{-4}$ for code and $2e^{-4}$ for math.
On both datasets, these best LoRA learning rates are underperformed by four alternative full finetuning learning rates. The best full finetuning learning rates are $5e^{-5}$ and $1e^{-5}$, respectively, an order of magnitude smaller than LoRA. For LoRA, we cannot find alternative learning rates that achieve at least 90% of the best learning rate’s performance. For full finetuning, there are two viable alternative learning rates for code and three for math.
Note that in these experiments, the LoRA models target all modules but the $W_{\text{gate}}$, with $\alpha=32$ which should preferably be higher for $r=64$. This explains the slight differences between Figures S1 and S3.
Figure S1: LoRA is more sensitive to learning rates compared to full finetuning. Llama-2-7B models (A) trained on Magicoder-Evol-Instruct-110k (Wei et al., 2023) and evaluated on HumanEval, (B) trained on MetaMathQA (Yu et al., 2023) and evaluated on GSM8K. Experiments here are performed with LionW; see Fig. S2 for a comparion to AdamW.
@@ B.1 Learning rate sensitivity analysis across optimizers
We compared the AdamW and Decoupled LionW optimizers by training for two epochs of Magicoder-Evol-Instruct-110K using different learning rates. We found that Decoupled LionW performed better on HumanEval for both LoRA and full finetuning, and across learning rates, as seen in Fig. S2.
@@
Figure S2: Comparing LionW to AdamW across learning rates for two epochs of the Magicoder-Evol-Instruct-110K dataset. Left: HumanEval; Right: Average of “Language Understanding” benchmarks in the MosaicML evaluation gauntlet. Both methods peak at the learning rate used in the original paper (Wei et al., 2023).
B.2 The importance of the alpha scaling parameter for LoRA
We found that the performance of all models was particularly sensitive to the LoRA $\alpha$ hyperparameter. Fig. S3 shows two experiments on two separate datasets (Magicoder-Evol-Instruct-110K and OpenWebMath) for LoRA with rank $r=256$. In both cases the best accuracy is achieved when $\alpha=2r$.
(a) Jointly sweeping over LoRA $\alpha$ and learning rate. The optimal choice is $\alpha=2r$ (blue).
(b) Continued pretraining with two different choices of $\alpha$, where $\alpha=2r$ is best (blue).
Figure S3: LoRA performance is sensitive to the $\alpha$ hyperparameter. We show that for Code IFT (a) and math CPT (b) an $\alpha$ that is scaled with rank such that $\alpha=2r$ leads to the highest accuracy.
Appendix C Finetuning on the Tülu-v2-mix dataset
We finetuned Llama-2-7B models on the Tülu-v2-mix (Ivison et al., 2023), which is a finetuning dataset containing chain of thought reasoning, multi-turn assistant conversations, math and science problems, code, and more. [^16] https://huggingface.co/datasets/allenai/tulu-v2-sft-mixture There are roughly 326k samples in this dataset.
As in all main experiments, we compared full finetuning and LoRA $r=16,64,256$, targeting all transformer modules. For each of the four experimental conditions, we trained a model for up to 6 epochs and evaluated it after 2, 4, and 6 epochs. Different from the main IFT experiments, the checkpoints evaluated are “hot” and are not cooled down for each training duration.
As in the original paper (Ivison et al., 2023), we assess math capabilities with GSM8K Cobbe et al. (2021), STEM, humanities, and social science capabilities as the average of 57 subjects of the Massive Multitask Language Understanding (MMLU; Hendrycks et al. (2020)), and conversational capabilities with Multi-Turn Benchmark (MT-bench (Zheng et al., 2024)) which includes 80 multi-turn conversations where the model responses are evaluated automatically by GPT-4. We also compute the same average forgetting score as in all other datasets in this paper.
Since datasets like Tülu-v2-mix are where LoRA is mostly used, we ask: can LoRA, even with a low rank, achieve full finetuning accuracy both in specific domains and in general conversational capabilities?
C.1 Experimental setup
After an initial learning rate sweep, we chose the following hyperparameters:
  • max_seq_len: 4096
  • optimizer: decoupled_lionw (betas=[0.9, 0.95])
  • learning_rate: Full finetuning: 5e-6; LoRA 1e-4
  • scheduler: cosine_with_warmup (alpha_f=0.01, t_warmup=0.1dur)
  • weight_decay: 0
  • precision: amp_bf16
  • global_train_batch_size: 192
  • device_train_microbatch_size: 6
  • gradient_clipping: norm (threshold=1)
  • num_gpus: 32
C.2 Results
First, we find that on MT-bench (Fig. S4), both LoRA and full finetuning meaningfully improve upon the base model (2.74), starting from the second epoch and improving only slightly when trained for longer. Crucially, all LoRA models are within one standard error of the mean of the full finetuning model (computed with 160 datapoints = 80 questions $\times$ 2 turns). That is, one can achieve full finetuning conversational capabilities with $r=16$. The caveat is that only 80 questions appear in this benchmark and that the variance, within model, is high.
In GSM8K (Fig. 5(a)), again, all models significantly improve upon the base model (0.145). Remarkably, even in this specific domain, LoRA and full finetuning are overlapping, with the best model being LoRA $r=256$ at epoch 4, which is followed by full finetuning at epoch 2. Here too, as in the other math datasets in the paper, there is an ordering by LoRA rank.
In MMLU (Fig. 5(b)), full finetuning and LoRA are overlapping with LoRA $r=64$ as the best model (epoch 4), followed by full finetuning at epoch 2. Here there is no ordering by rank.
As for forgetting (Fig. S6), we find an overall mild forgetting compared to the rest of the datasets in the paper. At two epochs, full finetuning does better than LoRA. The former starts to degrade at epoch 4. At epoch 6, the findings of the main paper are replicated: full finetuning forgets the most and we find a clear ordering of forgetting by rank.
Across all evaluations – learning and forgetting – full finetuning is the best model at epoch 2, and only degrades afterwards. LoRA, on the other hand, needs 4 epochs to train, mirroring the findings in the main part of the paper. LoRA $r=16$ seems to offer competitive conversational capabilities, and minimal forgetting, but it underperforms in domain-specific knowledge like math. Future work should seek to understand why this is the case.
Figure S4: Average MT-bench score with GPT-4 as a judge, calculated over 80 questions with two turns each. Base model value as reported in the MT-bench paper. We note that the Tülu paper reports a 6.3 MT-bench value from full finetuning of Llama-2-7B base model, which is only slightly exceeding the standard error from our average score.
(a) Accuracy in GSM8K.
(b) Average of MMLU benchmarks.
Figure S5: On Tülu-v2-mix, LoRA and full finetuning both improve upon the base model and perform comparably.
Figure S6: LoRA forgets less even on a more diverse IFT dataset like Tülu-v2-mix. We plot the average forgetting score, same as in all other datasets, as a function of training duration.
Appendix D Supplementary tables
Table S1: Starcoder-Python Results (HumanEval pass@1, temperature 0.2)
Num. tokens (billions) 0.25 0.50 1 2 4 8 16 20
Condition
LoRA (r=16) 0.143 0.144 0.141 0.141 0.154 0.159 0.162 0.162
LoRA (r=64) 0.142 0.146 0.141 0.153 0.157 0.176 0.194 0.196
LoRA (r=256) 0.144 0.142 0.143 0.159 0.159 0.208 0.211 0.224
Full Finetuning 0.152 0.153 0.172 0.181 0.218 0.258 0.255 0.263
Table S2: Starcoder-Python Results (Forgetting Average)
Num. tokens (billions) 0.25 0.50 1 2 4 8 16 20
Condition
LoRA (r=16) 0.645 0.642 0.645 0.642 0.644 0.640 0.638 0.635
LoRA (r=64) 0.646 0.644 0.646 0.646 0.639 0.634 0.626 0.626
LoRA (r=256) 0.644 0.645 0.643 0.639 0.636 0.630 0.618 0.617
Full Finetuning 0.625 0.624 0.625 0.616 0.599 0.583 0.551 0.545
Table S3: OpenWebMath Results (GSM8K)
Num. tokens (billions) 0.25 0.50 1 2 4 8 16 20
Condition
LoRA (r=16) 0.162 0.157 0.161 0.155 0.165 0.156 0.152 0.158
LoRA (r=64) 0.163 0.167 0.150 0.166 0.164 0.168 0.179 0.163
LoRA (r=256) 0.162 0.161 0.140 0.170 0.193 0.196 0.203 0.202
Full Finetuning 0.155 0.152 0.165 0.158 0.224 0.238 0.283 0.293
Table S4: OpenWebMath Results (Forgetting Average)
Num. tokens (billions) 0.25 0.50 1 2 4 8 16 20
Condition
LoRA (r=16) 0.640 0.641 0.646 0.641 0.643 0.641 0.636 0.637
LoRA (r=64) 0.640 0.640 0.638 0.637 0.643 0.634 0.634 0.627
LoRA (r=256) 0.638 0.638 0.637 0.634 0.633 0.620 0.620 0.616
Full Finetuning 0.634 0.634 0.640 0.630 0.629 0.619 0.613 0.618
Table S5: Magicoder-Evol-Instruct-110K Results (HumanEval pass@1)
Epoch 1 2 4 8 16
Condition
LoRA (r=16) 0.197 0.275 0.358 0.338 0.324
LoRA (r=64) 0.249 0.339 0.417 0.392 0.405
LoRA (r=256) 0.299 0.385 0.498 0.437 0.466
Full Finetuning 0.302 0.464 0.470 0.497 0.416
Table S6: Magicoder-Evol-Instruct-110K Results (Forgetting Average)
Epoch 1 2 4 8 16
Condition
LoRA (r=16) 0.653 0.648 0.652 0.646 0.609
LoRA (r=64) 0.652 0.651 0.632 0.580 0.510
LoRA (r=256) 0.655 0.659 0.631 0.552 0.517
Full Finetuning 0.595 0.579 0.512 0.446 0.414
Table S7: MetaMathQA Results (GSM8K)
Epoch 1 2 4 8 16
Condition
LoRA (r=16) 0.447 0.528 0.580 0.578 0.569
LoRA (r=64) 0.527 0.588 0.624 0.624 0.595
LoRA (r=256) 0.557 0.607 0.625 0.634 0.594
Full Finetuning 0.604 0.641 0.642 0.619 0.599
Table S8: MetaMathQA Results (Forgetting Average)
Epoch 1 2 4 8 16
Condition
LoRA (r=16) 0.628 0.617 0.616 0.616 0.596
LoRA (r=64) 0.617 0.609 0.608 0.586 0.568
LoRA (r=256) 0.613 0.607 0.599 0.584 0.567
Full Finetuning 0.598 0.599 0.590 0.572 0.559
Table S9: Tülu-v2-mix Results
Table S10: Tülu-v2-mix MT-Bench
Epoch 2 4 6
Condition
LoRA (r=16) 5.681 5.997 5.712
LoRA (r=64) 5.597 5.725 5.944
LoRA (r=256) 5.788 5.834 5.894
Full Finetuning 5.825 5.838 5.862
Table S11: Tülu-v2-mix MMLU
Epoch 2 4 6
Condition
LoRA (r=16) 0.491 0.502 0.504
LoRA (r=64) 0.503 0.509 0.504
LoRA (r=256) 0.502 0.496 0.492
Full Finetuning 0.507 0.504 0.502
Table S12: Tülu-v2-mix GSM8K
Epoch 2 4 6
Condition
LoRA (r=16) 0.251 0.275 0.280
LoRA (r=64) 0.285 0.270 0.295
LoRA (r=256) 0.296 0.335 0.301
Full Finetuning 0.324 0.291 0.303
Table S13: Tülu-v2-mix Forgetting Average
epoch 2 4 6
condition
LoRA (r=16) 0.650 0.657 0.657
LoRA (r=64) 0.649 0.655 0.647
LoRA (r=256) 0.653 0.649 0.629
Full Finetuning 0.660 0.652 0.621
Appendix E Supplementary Figures for SVD Analysis
@@
(a) Spectrum of $A$ and $A+cB$ as well as $cB$ for $c=0.1$. Notably, $A,cB,A+cB$ are all high rank.
@@
Figure S8: Analyzing the spectra of the sum of two $1000\times 1000$ Gaussian i.i.d matrices. $A$ and $B$ are $1000\times 1000$ random matrices with i.i.d. standard normal Gaussian entries.
Appendix F Solution Generation Diversity on HumanEval
For the best set of Llama-2-7B models trained on Magicoder for 4 epochs, we evaluate how the pass@$k$ metric in the HumanEval benchmark increases as we increase the parameter $k$ which controls the acceptance criterion. The pass@$k$ metric (Chen et al., 2021) is defined as
pass@$k$ $:=\mathop{\mathbb{E}}\left[1-\frac{{\binom{n-c}{k}}}{\binom{n}{k}}\right],$ (1)
where $n$ is the number of generations, $c$ the number of correct generations and $k$ determines the size of the sample set of generations considered for acceptance. Assuming we sample from the model outputs, i.e. sampling temperature $T>0$, then increasing $k$ will increase the diversity of generations, and increase the likelihood of a passing generation being present in a random subset of size $k$.
Figure S9 reports pass@$k$ for the LoRA models trained on the Magicoder dataset as well as the base Llama-2-7B model. For all models, as we increase $k$, the pass@$k$ consistently and monotonically improves. Finetuned models scores are substantially higher than the base model. At $k=1$, full finetuning outperforms the LoRA model whose scores are ordered from largest to smallest rank, as expected. As $k$ increases we observe that all models improve their pass@$k$ scores, and that the gap between them reduces when $k>16$. We note that full finetuning is superior across all values of $k$ with temperature 0.8. This complements the results in Fig. 1 which used a temperature of 0.2 and pass@$1$, where the improvements upon $r=256$ at epoch 4 are less clear.
Figure S9: HumanEval pass@$k$ for models trained on the Magicoder dataset. For every model, we sample 256 independent generations with temperature 0.8.
Appendix G Training Datasets
G.1 MetaMathQA (Math IFT)